# AGRAG: Advanced Graph-based Retrieval-Augmented Generation for LLMs

### v1

@@ III-B Step 1: Data Preparation
In this subsection, we utilize a TFIDF [37] based entity extraction approach to substitute the LLM-based approach of other Graph-based RAG models, reduce LLM hallucination, resulting in less noise generated from the early entity extraction step, and hence less error propagation, address the inaccurate graph construction issue, lead to better effectiveness. The pseudo code of AGRAG’s graph construction procedure is shown in algorithm 1.
@@ 1 Let $\mathcal{V}\leftarrow\emptyset$, $\mathcal{E}\leftarrow\emptyset$;
4   Extract entity set $\mathcal{V}_{t}$ based on Equation 3;
@@ 12         Add passage edge between $v$ and $t$: $\mathcal{E}\leftarrow\mathcal{E}\cup\{e_{(v,t)}\}$;
To begin with, we split texts in $\mathcal{T}$ into fix lengthed chunks with a certain amount of token overlap between chunks:
@@
where $l_{t}$ and $l_{o}$ denotes the length of text chunks and token overlap between chunks. Next, we apply the statistics-based entity extraction approach on the text chunks $\mathcal{T}_{c}$. Specifically, our approach modifies the TFIDF score [37] as the entity extraction score $\text{ER}(v,t)$ for each term $v$ in text chunk $t$ as an entity, to make the value of ER fall in $(0,1)$. Here, our entity detection is based on an n-gram setting with the maximum n-gram (number of words in an entity) as $b$, and the equation of $\text{ER}(v,t)$ is as follow:
@@
where $\mathcal{T}_{c}$ denotes all text chunks obtained from Equation 1, $\text{count}(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}_{c}|$ denotes the number of texts in the corpus $\mathcal{T}_{c}$ that contain the keyword $v$. Here, our entity extraction procedure will extract words or phrases with $\text{ER}(v,t)>\tau$ as the entity nodes $\mathcal{V}_{t}$ for text chunk $t$:
@@
Entities in $\mathcal{V}_{t}$ are later served as graph nodes, hence we denote them as entity nodes. For our entity extraction procedure, the entity boundary $b$ and entity extraction threshold $\tau$ are considered as hyper-parameters, we further analyze their setting in subsection IV-E.
After entity extraction, we extract relations from each text chunk $t\in\mathcal{T}_{c}$ based on its extracted entity $\mathcal{V}_{t}$. Similar from other Graph-based RAG models, we also apply LLM at the relation extraction step:
@@
where $\mathcal{V}_{t}$ denotes the extracted entities of text chunk $t$, and in the relation extraction prompt $p_{\text{re}}$, we ask the LLM to detect relations between entities in $\mathcal{V}_{t}$ w.r.t. text chunk $t$. Although we also apply LLM for relation extraction, we input less noise in this step, mitigating the error propagation, as our statistics-based entity extraction avoids LLM hallucination, hence resulting in a more comprehensive and less noisy $\mathcal{V}_{t}$ as the Equation 4 input.
After relation extraction, we add text chunks in $\mathcal{T}_{c}$ to $\mathcal{V}$ as passage nodes, and add passage edges $e_{(v,t)}$ between entities the text chunks contain them to $\mathcal{E}$, the respective relation of these edges are namely Contains. By introducing passage nodes, we integerate contextual information into the constructed KG, and the entity-passage edges can capture the inter-text correlations, make our model able to generate reasoning paths across different text chunks [31].
Furthermore, we add synonym edges $\mathcal{E}_{\text{syn}}$ between entity nodes in $\mathcal{V}$ that have similarties larger than the synonym threshold $\phi$, the relation of these edges are namely synonym:
@@
here, if the entity $v$ not exist in text chunk $t$, $e_{(v,t)}=\emptyset$.
Figure 3: An example of the MCMI generated. Red triples indicate those mapped from the query. Our algorithm first constructs the MCMI based on a Steiner tree rooted at these mapped triples, then iteratively expands it by adding neighboring triples with the highest influence-cost ratio $s/c$, continuing until no neighboring triples have this ratio greater than the average influence-cost ratio of the current graph, covering more query related triples (in orange).
@@ III-C Step 2: Graph Retrieval
In this subsection, we introduce the graph retrieval procedure of AGRAG. Given the KG $\mathcal{KG}(\mathcal{V},\mathcal{E})$ constructed at the Step 1, we first map the query $q$ to $\mathcal{KG}$ and obtain raw mapped triplet facts $\mathcal{F}_{\text{raw}}$. Then we encode each triplet fact in KG $f\in\mathcal{F}_{\mathcal{KG}}$ and the query $q$ into vector representations q and f with encoder $\mathcal{M}$: $\textbf{q}\leftarrow\mathcal{M}(q)$, then select facts with top-k dense semantic mapping score MS to query $q$ based on their cosine similarity:
@@
where $k_{a}$ is a hyperparameter that controls the number of raw mapped triple extracted, and f are obtained through encoding the text concatenation of entities and relation of the triplet fact.
Further, we apply LLM based filter on raw mapped triplet facts $\mathcal{F}_{\text{raw}}$ to generate the mapped triplet facts $\mathcal{F}$, to reduce the noise in mapped triples:
@@
Next, based on the mapped triplet facts $\mathcal{F}$ and the KG structure, we calculate the Personalized PageRank (PPR) [33] score as our node influence score $\mathcal{S}_{\mathcal{V}}$ for node set $\mathcal{V}$:
$\mathcal{S}_{\mathcal{V}}=\lim_{i\to\infty}\text{PPR}^{i},\,\text{PPR}^{i}=(1-d)\times\mathbf{p}+d\times\mathbf{P}\cdot\text{PPR}^{i-1}.$ (10)
To align with [31], we set the damping factor $d$ to 0.5, and converge the PPR calculation when it changes less than $10^{-7}$; $\mathbf{p}\in\mathbb{R}^{|\mathcal{V}|}$ denotes the personalization vector that depends on the mapped triples $\mathcal{F}$. The $\mathbf{P}\in\mathbb{R}^{|\mathcal{V}\times\mathcal{V}|}$ in Equation 10 denotes the transition probability matrix, and can be obtained by:
$\mathbf{P}_{u,v}=\begin{cases}\frac{1}{|\mathcal{N}(u)|},&\text{if }e_{(u,v)}\in\mathcal{E},\\ \frac{1}{|\mathcal{V}|},&\text{otherwise},\end{cases}$ (11)
where $\mathcal{N}(u)$ denotes the 1-hop neighbor set of node $u$ in the KG. Moreover, following [31], after calculate the PPR score, we multiply the PPR score of all passage nodes by a balance factor 0.05, to avoid over concentration on them. After that, we can obtain the node influence score $s_{v}$ for each node $v\in\mathcal{V}$ as follow:
$s_{v}=\mathcal{S}_{\mathcal{V}}(v),\quad\forall v\in\mathcal{V}.$ (12)
@@
1 Assign node weights $\mathcal{S}_{\mathcal{V}}$ with PPR score w.r.t. $\mathcal{F}$ based on Equation 10: $\mathcal{S}_{\mathcal{V}}\leftarrow\lim_{i\to\infty}\text{PPR}^{i}$;
@@
3   Assign edge costs $c_{e_{(u,v)}}$ based on Equation 13: $c_{e_{(u,v)}}\leftarrow(1-\text{MS}(\mathbf{q},\mathbf{f}_{e_{(u,v)}}))/2$;
@@ 4 $\mathcal{C}_{\mathcal{E}}\leftarrow\mathcal{C}_{\mathcal{E}}\cup\{c_{e_{(u,v)}}\}$;
At last, we calculate the edge cost $\mathcal{C}_{\mathcal{E}}$ based on the similarity between query $q$’s representation q and relation edge $e$’s respective triplet facts $f_{e}$’s representation $\textbf{f}_{e}$:
$\mathcal{C}_{\mathcal{E}}=\{c_{e}\mid e\in\mathcal{E}\},\quad c_{e}=\frac{1-\text{MS}(\mathbf{q},\mathbf{f}_{e})}{2}.$ (13)
Moreover, to guarantee the KG’s connectivity, we connect all passage nodes to a pseudo node, with the pseudo edge (relation) namely pseudo_relation, and the cost of these pseudo edges is set to 10 to penalize paths through the pseudo node, aligning with the setting of HippoRAG2 [31].
@@
Given an node and edge weighted connected graph $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$, a set of mapped triples $\mathcal{F}$. The aim of this problem is to generate a subgraph $\mathcal{G}_{\text{MCMI}}$ with the maximum average node score and minimum total edge cost that contains all the mapped triples.
Next, we prove our MCMI Subgraph Generation problem can be reduced from the NP-hard Steiner Tree problem [70].
@@
The Minimum Cost Maximum Influence Subgraph Generation problem is NP-hard.
Proof.
We define the decision version of MCMI as follows: Given a weighted knowledge graph $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$, entity node set from mapped triples (considered as terminal nodes) $\mathcal{V}_{\text{term}}\subseteq\mathcal{V}$, and two thresholds $\alpha\in(0,1]$ and $\beta\in\mathbb{R}^{+}$, does there exist a connected subgraph $\mathcal{G}^{\prime}(\mathcal{V}^{\prime},\mathcal{E}^{\prime})\subseteq\mathcal{KG}$ such that:
  1. (i)
    $\mathcal{V}_{\text{term}}\subseteq\mathcal{V}^{\prime}$, which means all mapped triples are included,
  2. (ii)
    the average node influence score satisfies $\frac{1}{|\mathcal{V}^{\prime}|}\sum_{v\in\mathcal{V}^{\prime}}s_{v}\geq\alpha$, and
  3. (iii)
    the total edge cost satisfies $\sum_{e\in\mathcal{E}^{\prime}}c_{e}\leq\beta$?
Now, consider an arbitrary MCST instance with graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, terminal node set $\mathcal{V}_{\text{term}}\subseteq\mathcal{V}$, and total cost $\gamma$. We construct an MCMI instance as follows:
  1. 1.
    Use the same graph $\mathcal{G}$ as the KG $\mathcal{KG}$.
  2. 2.
    Set $s_{v}=1$ for all $v\in\mathcal{V}$, hence any subgraph has average node score exactly 1.
  3. 3.
    Keep edge costs $c_{e}$ unchanged for each $e\in\mathcal{E}$.
  4. 4.
    Set $\alpha=1$ and $\beta=\gamma$.
Because all nodes have score 1, condition (ii) holds for any non-empty subgraph. Therefore, a Steiner tree of cost $\leq\gamma$ exists in the MCST instance if and only if a feasible subgraph exists in the MCMI instance. This establishes a polynomial-time reduction, proving that MCMI is NP-hard. ∎
Since the MCMI Subgraph Generation problem is the NP-hard, it is infeasible to obtain the optimal result in polynomial time. Therefore, we propose a greedy algorithm (algorithm 3) based on the Mehlhorn’s algorithm [71] to solve this problem.
To begin with, we initialize the MCMI as the Minimum Cost Steiner Tree (MCST) through all mapped triplet facts $\mathcal{F}$ in the weighted KG $\mathcal{KG}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$ (line 1 to line 10 of algorithm 3), where score $s_{v}\in\mathcal{S}_{\mathcal{V}}$ of each node $v$, and cost $c_{e}\in\mathcal{C}_{\mathcal{E}}$ of each edge $e$ are obtained with Equation 12 and Equation 13, respectively:
$\mathcal{G}_{\text{MCMI}}(\mathcal{V}_{\text{MCMI}},\mathcal{E}_{\text{MCMI}})=\text{MCST}(\mathcal{F},\mathcal{KG}),$ (14)
where $\text{MCST}(\mathcal{F},\mathcal{KG})$ denotes the Minimum Cost Steiner tree generation algorithm.
Next, we calculate the ratio of the nodes’ influence score and their connected edges’ cost (i.e., $s/c$), and check this ratio of all nodes in the current MCMI’s 1-hop neighborhood $\mathcal{N}(\mathcal{G}_{\text{MCMI}})$, if this ratio is greater than that the current MCMI subgraph’s average score-cost ratio, we add this node (entity) and the corresponding edge (relation) to the MCMI, recalculate the MCMI’s score-cost ratio and continue the iteration, until all neighbor node’s score-ratio is less or equal than the MCMI subgraph’s average score-cost ratio. That is, for each edge $e_{(u,v)}$, we have (line 10 to line 16 of algorithm 3):
@@
where $r_{\text{MCMI}}$ denotes the average score-cost ratio of the current MCMI subgraph; $\mathcal{V}_{\text{MCMI}}$ and $\mathcal{E}_{\text{MCMI}}$ denotes the node and edge set of current MCMI subgraph; $e_{(\text{MCMI},v)}$ denotes the outgoing edges from the MCMI subgraph to its 1-hop neighborhood; And $r_{\text{MCMI}}$ is calculated as follow:
$r_{\text{MCMI}}=\sum_{e_{(u,v)}\in\mathcal{E}_{\text{MCMI}}}\frac{s_{u}+s_{v}}{2\times c_{e_{(u,v)}}\times|\mathcal{E}_{\text{MCMI}}|}.$ (16)
After generate MCMI subgraph generation, we serialize the generated MCMI subgraph into a text graph string $g$:
@@
Running Example: As shown in Figure 3, if we simply calculate the minimum cost Steiner tree within the given KG, the algorithm will simply return the two mapped fact (Kidney, Work, Kidney cancer), and (Kidney, Spread, Tumors) as the reasoning path, and the LLM’s response answer would likely to be Tumors. However, if we go one step further, then within the generated cyclic graph, we can have more complete reasoning information regarding the query, and involve the correct answer Blood vessels.
@@
The approximation ratio of the MCMI Subgraph Generation Algorithm (algorithm 3) is 2.
Proof.
In algorithm 3, we firstly compute the minimum cost Steiner tree $\mathcal{G}(\mathcal{V}_{\text{MCST}},\mathcal{E}_{\text{MCST}})$ which connects all nodes in $\mathcal{V}_{\text{term}}$. Suppose the total cost for $\mathcal{G}$ is $\gamma$, and the optimal minimum cost Steiner tree is $\mathcal{G}^{*}(\mathcal{V}_{\text{MCST}^{*}},\mathcal{E}_{\text{MCST}^{*}})$ with cost $\gamma^{*}$. $\mathcal{G}^{*}$ covers at least all edges from $\mathcal{G}$. Following triangle inequality, for $\forall e\in\mathcal{E}_{\text{MCST}^{*}}$, the sum of weight is less than 2 times of edge weights in $\mathcal{E}_{\text{MCST}}$, which is $\gamma\leq 2\gamma^{*}$. The process from line 12 to line 15 will retrieve more edges. However, they will not change the weight in $\mathcal{G}$. Hence, the approximation ratio of algorithm 3 is 2. ∎
@@
1 Compute a minimum cost Steiner tree MCST $\mathcal{G}$ of $\mathcal{KG}$;

### v2

@@ III-B Step 1: Data Preparation
In this subsection, we utilize a TFIDF [64] based entity extraction approach to substitute the LLM-based approach of other Graph-based RAG models, reduce LLM hallucination, resulting in less noise generated from the early entity extraction step, and hence less error propagation, address the inaccurate graph construction issue, leading to better effectiveness. The pseudo code of AGRAG’s graph construction procedure is shown in Algorithm 1.
@@ 1 Let $\mathcal{V}\leftarrow\emptyset$, $\mathcal{E}\leftarrow\emptyset$;
4   Extract entity set $\mathcal{V}_{t}$ based on Equation (3);
@@ 12         Add passage edge between $v$ and $t$: $\mathcal{E}\leftarrow\mathcal{E}\cup\{e_{(v,t)}\}$;
To begin with, we split texts in $\mathcal{T}$ into fixed-length chunks with a certain amount of token overlap between chunks:
@@
where $l_{t}$ and $l_{o}$ denotes the length of text chunks and token overlap between chunks. Next, we apply the statistics-based entity extraction approach on the text chunks $\mathcal{T}_{c}$. Specifically, our approach modifies the TFIDF score [64] as the entity extraction score $\text{ER}(v,t)$ for each term $v$ in text chunk $t$ as an entity, to make the value of ER fall in $(0,1)$. Here, our entity detection is based on an n-gram setting with the maximum n-gram (number of words in an entity) as $b$, and the equation of $\text{ER}(v,t)$ is as follows:
@@
where $\mathcal{T}_{c}$ denotes all text chunks obtained from Equation (1), $\text{count}(v,t)$ is the number of times that the term $v$ appears in the text $t$, and $|t_{j}:v\in t_{j},\ t_{j}\in\mathcal{T}_{c}|$ denotes the number of texts in the corpus $\mathcal{T}_{c}$ that contain the keyword $v$. Here, our entity extraction procedure will extract words or phrases with $\text{ER}(v,t)>\tau$ as the entity nodes $\mathcal{V}_{t}$ for text chunk $t$:
@@
Entities in $\mathcal{V}_{t}$ are later served as graph nodes, hence we denote them as entity nodes. For our entity extraction procedure, the entity boundary $b$ and entity extraction threshold $\tau$ are considered as hyper-parameters, we further analyze their setting in Section IV-E.
After entity extraction, we extract relations from text chunks $t\in\mathcal{T}_{c}$ based on its extracted entities $\mathcal{V}_{t}$. Align with other Graph-based RAG models, we also apply LLM at this step:
@@
where $\mathcal{V}_{t}$ denotes the extracted entities of text chunk $t$, and in the relation extraction prompt $p_{\text{re}}$, we ask the LLM to detect relations between entities in $\mathcal{V}_{t}$ w.r.t. text chunk $t$. Although we also apply LLM for relation extraction, we input less noise in this step, mitigating the error propagation, as our statistics-based entity extraction avoids LLM hallucination, hence resulting in a more comprehensive and less noisy $\mathcal{V}_{t}$ as the Equation (4) input.
After relation extraction, we add text chunks in $\mathcal{T}_{c}$ to $\mathcal{V}$ as passage nodes, and add passage edges $e_{(v,t)}$ between entities contained within the text chunks to $\mathcal{E}$, the respective relation of these edges is namely Contains. By introducing passage nodes, we integrate contextual information into the constructed KG, and the entity-passage edges can capture the inter-text correlations, make our model able to generate reasoning paths across different text chunks [gutiérrez2025ragmemory].
Furthermore, we add synonym edges $\mathcal{E}_{\text{syn}}$ between entity nodes in $\mathcal{V}$ that have similarities larger than the synonym threshold $\phi$, the relation of these edges is namely synonym:
@@
here, if the entity $v$ does not exist in text chunk $t$, $e_{(v,t)}=\emptyset$.
Figure 2: An example of MCMI. Red triples indicate those mapped from the query. Our algorithm first constructs the MCMI based on a Steiner tree rooted at these mapped triples, then iteratively expands it by adding neighboring triples (in orange) with the lowest cost-score ratio, continuing until no neighboring triple has its ratio less than the current MCMI’s ratio.
@@ III-C Step 2: Graph Retrieval
In this subsection, we introduce the graph retrieval procedure of AGRAG. Given the KG $\mathcal{KG}(\mathcal{V},\mathcal{E})$ constructed at the Step 1, we first map the query $q$ to $\mathcal{KG}$ and obtain raw mapped triplet facts $\mathcal{F}_{\text{raw}}$. To achieve this, we encode each triplet fact in KG $f\in\mathcal{F}_{\mathcal{KG}}$ and the query $q$ into vector representations q and f with encoder $\mathcal{M}$: $\textbf{q}\leftarrow\mathcal{M}(q)$, then select facts with top-k dense semantic mapping score MS to query $q$ based on their cosine similarity:
@@
where $k_{a}$ is a hyperparameter that controls the number of raw mapped triples extracted, and f are obtained through encoding the text concatenation of entities and relation of the triplet fact.
Further, as the triplets in KG are extracted in a query unaware manner, query irrelevant triplets can exist in the raw mapped triplet facts $\mathcal{F}_{\text{raw}}$, hence we apply LLM-based filter on $\mathcal{F}_{\text{raw}}$ to generate the mapped triplet facts $\mathcal{F}$, to reduce the query irrelevant noise:
@@
Next, based on the mapped triplet facts $\mathcal{F}$ and the KG structure, we calculate the Personalized PageRank (PPR) [27] score as our node influence score $\mathbf{s}$ for node set $\mathcal{V}$:
$\mathbf{s}=\lim_{i\to\infty}\text{PPR}^{i},\,\text{PPR}^{i}=(1-d)\times\mathbf{p}+d\times\mathbf{P}\cdot\text{PPR}^{i-1},$ (10)
and we have the node score set $\mathcal{S}_{\mathcal{V}}=\{\mathbf{s}[v]|v\in\mathcal{V}\}$. To align with [gutiérrez2025ragmemory], we set the damping factor $d$ to 0.5, and converge the PPR calculation when it changes less than $10^{-7}$; $\mathbf{p}\in\mathbb{R}^{|\mathcal{V}|}$ denotes the personalization vector that depends on the mapped triples $\mathcal{F}$:
$\mathbf{p}[u]=\begin{cases}1,&\text{if }u\in\mathcal{F},\\ 0,&\text{otherwise}.\end{cases}$ (11)
The $\mathbf{P}\in\mathbb{R}^{|\mathcal{V}\times\mathcal{V}|}$ in Equation (10) denotes the transition probability matrix, and can be obtained by:
$\mathbf{P}[u][v]=\begin{cases}\frac{1}{|\mathcal{N}(u)|},&\text{if }e_{(u,v)}\in\mathcal{E},\\ \frac{1}{|\mathcal{V}|},&\text{otherwise},\end{cases}$ (12)
where $\mathcal{N}(u)$ denotes the 1-hop neighbor set of node $u$ in the KG. To ensure the PPR’s faster convergence and more stable iterative computation [7], as well as normalizing the score distribution [4]. Following [gutiérrez2025ragmemory], we set the transition probability for non-edges to $\frac{1}{|\mathcal{V}|}$, and after calculating the PPR score, we multiply the PPR score of all passage nodes by a balance factor 0.05, to avoid over concentration on them. After that, we can obtain the node influence score $s_{v}$ for each node $v\in\mathcal{V}$ as follows:
$s_{v}=\mathbf{s}[v],\quad\forall v\in\mathcal{V}.$ (13)
@@
1 Assign node weights $\mathcal{S}_{\mathcal{V}}$ with PPR score w.r.t. $\mathcal{F}$ based on Equation (10): $\mathbf{s}\leftarrow\lim_{i\to\infty}\text{PPR}^{i}$;
@@
3   Assign edge costs $c_{e_{(u,v)}}$ based on Equation (14): $c_{e_{(u,v)}}\leftarrow(1-\text{MS}(\mathbf{q},\mathbf{f}_{e_{(u,v)}}))/2$;
@@ 4 $\mathcal{C}_{\mathcal{E}}\leftarrow\mathcal{C}_{\mathcal{E}}\cup\{c_{e_{(u,v)}}\}$;
At last, we calculate the edge cost $\mathcal{C}_{\mathcal{E}}$ based on the similarity between query $q$’s representation q and relation edge $e_{(u,v)}$’s respective triplet facts $f_{e_{(u,v)}}$’s representation $\textbf{f}_{e_{(u,v)}}$:
$\mathcal{C}_{\mathcal{E}}=\{c_{e_{(u,v)}}\mid e_{(u,v)}\in\mathcal{E}\},\quad c_{e_{(u,v)}}=\frac{1-\text{MS}(\mathbf{q},\mathbf{f}_{e_{(u,v)}})}{2}.$ (14)
Moreover, to guarantee the KG’s connectivity, we connect all passage nodes to a pseudo node, with the pseudo edge (relation) namely pseudo_relation, and the cost of these pseudo edges is set to 10 to penalize paths through the pseudo node, aligning with the setting of HippoRAG2 [gutiérrez2025ragmemory].
@@
Given a node-weighted and edge-weighted, connected graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$ and a set of mapped triples $\mathcal{F}$, let $\mathcal{V}_{\text{term}}$ be the set of nodes in $\mathcal{F}$, called terminals. Let $r\in\mathcal{V}_{\text{term}}$ be a designated root terminal. The goal is to generate a subgraph $\mathcal{G}_{\text{MCMI}}(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{S}^{\prime}_{\mathcal{V}},\mathcal{C}^{\prime}_{\mathcal{E}})$ that contains all terminals $\mathcal{V}_{\text{term}}$ such that every terminal $v\in\mathcal{V}_{\text{term}}$ is reachable from root terminal $r$, while minimizing the total cost-score ratio $r(\mathcal{G_{\text{MCMI}}})$:
$r(\mathcal{G}_{\text{MCMI}})=\sum_{e_{(u,v)}\in\mathcal{E}^{\prime}}\frac{c_{e_{(u,v)}}}{s_{u}+s_{v}},$ (15)
where $s_{v}=\mathbf{s}[v]$ denotes the Personalized PageRank (PPR) score of node $v$, as defined in Equation (10), and $c_{e_{(u,v)}}$ represents the edge cost of $e_{(u,v)}$, as defined in Equation (14).
By minimizing the cost-score ratio in Equation (15), we can ensure the generated subgraph maximizes the inclusion of high-influence, query-relevant nodes while strictly penalizing low-similarity connections to effectively filter out noise.
@@
The MCMI subgraph generation is NP-hard.
Proof Sketch.
We establish NP-hardness by constructing a bijection between the MCMI problem and the classic Steiner Tree problem via a transformed weight function. Due to the space limitation, we put our proof at https://github.com/Wyb0627/AGRAG/blob/main/proof.pdf. ∎
Since the MCMI Subgraph Generation problem is NP-hard, it is infeasible to obtain the optimal result in polynomial time. Therefore, we propose a greedy algorithm (Algorithm 3) based on the Mehlhorn’s algorithm [50] to solve this problem.
To begin with, we initialize the MCMI as the Minimum Cost Steiner Tree (MCST) through all mapped triplet facts $\mathcal{F}$ in the weighted graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathcal{S}_{\mathcal{V}},\mathcal{C}_{\mathcal{E}})$ (line 1 to line 10 of Algorithm 3), where score $s_{v}\in\mathcal{S}_{\mathcal{V}}$ of each node $v$, and cost $c_{e_{(u,v)}}\in\mathcal{C}_{\mathcal{E}}$ of each edge $e$ are obtained with Equation (13) and Equation (14), respectively:
$\mathcal{G}_{\text{MCMI}}(\mathcal{V}_{\text{MCMI}},\mathcal{E}_{\text{MCMI}})=\text{MCST}(\mathcal{F},\mathcal{G}),$ (16)
where $\text{MCST}(\mathcal{F},\mathcal{G})$ denotes the Minimum Cost Steiner tree generation algorithm.
Next, we calculate the ratio of the edges’ cost and their connected nodes’ influence score (i.e., $c/s$), and check this ratio of all nodes in the current MCMI’s 1-hop neighborhood $\mathcal{N}(\mathcal{G}_{\text{MCMI}})$, if this ratio is less than that of the current MCMI subgraph’s total cost-score ratio, we add this node (entity) and the corresponding edge (relation) to the MCMI, recalculate the MCMI’s cost-score ratio and continue the iteration, until all neighbor node’s cost-score ratio is more than or equal to the MCMI subgraph’s total cost-score ratio. That is, for each edge $e_{(u,v)}$, we have (line 10 to line 16 of Algorithm 3):
@@
where $\mathcal{V}_{\text{MCMI}}$ and $\mathcal{E}_{\text{MCMI}}$ denote the node and edge set of current MCMI subgraph; $e_{(\text{MCMI},v)}$ denotes the outgoing edges from the MCMI subgraph to its 1-hop neighborhood.
After the MCMI subgraph generation, we serialize the generated MCMI subgraph into a text graph string $g$:
@@
Running Example: As shown in Figure 2, if we simply calculate the minimum cost Steiner tree within the given KG, the algorithm will simply return the two mapped fact (Kidney, Work, Kidney cancer), and (Kidney, Spread, Tumors) as the reasoning path, and the LLM’s response answer would likely be Tumors. However, if we go one step further, then within the generated cyclic graph, we can have more complete reasoning information regarding the query, and involve the correct answer Blood vessels.
We further prove our MCMI generation algorithm’s MCST generation step (Lines 1-10 in Algorithm 3) can still have a 2-approximation rate in Theorem 2. And its MCMI generation step (Lines 11-15 in Algorithm 3) strictly monotonically gains information and terminates within at most $|\Omega|$ iterations in Theorem 3, where $\Omega$ denotes the candidate edge set of MCMI. Together, they establish the effectiveness and efficiency of our MCMI generation Algorithm 3.
@@
The MCST generation step (Lines 1-10 in Algorithm 3) guarantees a 2-approximation.
Proof Sketch.
We bound the cost of the generated subgraph by comparing the Minimum Spanning Tree (MST) on the auxiliary graph $H$ to a shortcutted Euler tour (by depth-first traversal) of the optimal Steiner tree. Due to space limitation, we put our proof at https://github.com/Wyb0627/AGRAG/blob/main/proof.pdf. ∎
Theorem 3.
The greedy expansion of the MCMI generation step (Lines 11-15 in Algorithm 3) strictly monotonically gains information and terminates within at most $|\Omega|$ iterations, where $\Omega$ denotes the candidate edge set of MCMI:
$\Omega=\left\{e_{(u,v)}\in\mathcal{E}\setminus\mathcal{E}_{\text{MCST}}\mid\frac{c_{e_{(u,v)}}}{s_{v}}<r(\mathcal{G}_{\text{MCST}})\right\}$ (19)
Proof Sketch.
We prove the MCMI generation steps’ strict monotonic improvement by showing the algorithm only adds edges if their marginal cost-score ratio is strictly lower than the current subgraph’s ratio. Termination is guaranteed because the expansion process is restricted to a finite set of candidate edges. Due to the space limitation, we put our proof at https://github.com/Wyb0627/AGRAG/blob/main/proof.pdf. ∎
@@
1 Compute an approximated minimum cost Steiner tree (MCST) $\mathcal{G}$ of $\mathcal{KG}$;